OpenStudy (anonymous):

a screwdriver is dropped from the top of an elevator shaft. exactly 4.5 sec later, the sound of the scrwedriver hitting bottom is heard. how deep is the shaft? the distance that a dropped object falls in t sec is represented by the formula s=16t^2 the speed of sound is 1100 ft/sec

OpenStudy (anonymous):

use s = 16t^2 to find the height s = 16 * (4.5) ^2 = 324 ft now this distance is calculated with the extra time (the time taken for the sound to travel from bottom to top) so using the distnce u calculated, fiind the time taken for the sound to reach up s = ut 324 = 1100 * t t = 0.294s now subtract this time from the whole time (4.5) and substitute in the equation s = 16t^2 s = 16* (4.5-0.294)6^2 s = ?

OpenStudy (anonymous):

DID U GET IT?